# Multilevel Sampling in Algebraic Statistics

### v1

@@ 1 Introduction
Henry Wynn’s work in bringing modern algebraic geometry to the field of statistics laid the foundation for the development of algebraic statistics as a subject. The term ‘algebraic statistics’ was coined in his joint work [33] (see also the monograph [32]) and broadly denotes the application of nonlinear algebraic and nonconvex geometric methods to statistical problems. Although algebraic and geometric ideas had long been present in statistics, the predominant tools before the 2000s were linear algebra and convex geometry. A shift began in the mid-1990s: Pistone and Wynn [34] employed Gröbner bases to study confounding in experimental design, while Diaconis and Sturmfels [12] used them for performing exact conditional tests. Wynn and collaborators went further, recasting core concepts in statistics–such as optimal design and system reliability theory [30]–within the framework of computational algebra. Today, Wynn’s influence remains central in these areas. A central theme of his work was bridging interdisciplinary language barriers to connect ideas across fields. This paper presents one such effort [^1] In 2016, Henry Wynn visited Chicago twice. During discussions with S. Petrović, F. Hickernell, and their students, he proposed a new approach to sampling lattice points in polytopes, inspired by multilevel methods.: the adaptation of multilevel sampling [17]–a powerful tool in Monte Carlo methods employed to reduce computational cost when approximating expectations of quantities of interest, most commonly applied to path simulation [16]–to the setting of sampling discrete structures, a core concern in algebraic statistics.
@@
Sampling algorithms and strategies for exploring the set (1) abound. These include, for example, the hit-and-run sampler [40] (see also [4]), sequential importance sampling (SIS, see [29] as a standard reference), and a hybrid approach that combines SIS and Markov bases [27]. [9] reviews and compares a Markov chain implementation and SIS for logistic regression models, showing how SIS outperforms in certain scenarios; while [14] demonstrates the superiority of Markov bases over SIS in general fibers. Dobra also focused on the problem of actually computing Markov bases in practice, which is intractable for many problems of reasonable size that arise in applications. Indeed, there have been many attempts to work without the entire basis; see [22] for an early idea based on remarks made in [12]. The fact of the matter is that using a proper subset of a Markov basis necessarily disconnects some fibers, and it is impossible to tell in general which ones. A recent approach to circumvent the problem is a Bayesian framework, called RUMBA [6], that learns combinations of smaller basis elements. Given recent advances in reinforcement learning, another reasonable way to try to solve the problem is by actor-critic algorithms; this has been proposed in [20], motivated by [38]. Today, the fiber sampling problem still faces some open challenges: training agents comes at a cost, tuning sampler parameters takes time, and mixing of Markov chains is difficult to evaluate in a timely manner.
@@ 1.1 Problem and notation
Most of our readers will be familiar with the fact that conditional distributions given sufficient statistics in log-affine models, as defined in [28], are supported exactly on lattice points in polytopes. Log-affine models are exponential family models ([7]) for $k$ discrete random variables $X_{1},\dots,X_{k}$ with finite state spaces, $X_{i}\subset[d_{i}]:=\{1,\dots,d_{i}\}$. An exponential family being log-affine implies the existence of a large, often sparse matrix $A=[a_{1}\dots a_{n}]\subset\mathbb{Z}^{d\times n}$, which computes the model’s sufficient statistics: the family of probability distributions $\mathcal{M}_{A}:=\{p_{\theta}\}_{\theta\in\Theta}$ takes the form:
@@
One canonical application for constructing a sample from (1) is to perform the test of goodness-of-fit for $\mathcal{M}_{A}$ when asymptotic considerations fail–namely, in high-dimensional or large but possibly sparse state spaces (see [21] for an early discussion of this issue). Thus, in solving the problem, we focus on the chi-square statistic, as it can be used as a goodness-of-fit statistic for testing the fit of the model to the observed data. The distribution on the fiber can be either uniform (for conditional volume tests [11]) or the conditional distribution given sufficient statistics (uniform or often hypergeometric, depending on the table cell sampling scheme [15]). Measuring distance of observed data to the set of joint observations consistent with the log-linear model naturally lead to conditioning on sufficient statistics, removing dependence on model parameters; conditional inference was described by Agresti in 1990 (see [2, 1]). This was the context considered in [12], who introduced Markov bases for solving this broad class of sampling problems. The crucial fact is that a Markov chain is built using moves that have analogues in polynomial algebra and polyhedral geometry. Each instance inherits a finite set of moves necessary to obtain an irreducible chain for any value of the sufficient statistic. Diaconis and Sturmfels named any such finite collection of moves a Markov basis; see [12, Theorem 3.1]. Their seminal paper has over 1000 citations; we point the readers to the comprehensive references [5] and [36] and a recent survey [3].
@@
To fix the notation, we will refer to the set ${\mathcal{F}_{A}(b)}$ from Equation (1) as the $b$-fiber of $A$. Elements ${\mathcal{B}}\subset\ker_{\mathbb{Z}}(A)$, which by definition preserve the value of the sufficient statistics, are called moves on the fiber. For given moves $\mathcal{B}$, one can construct a fiber graph with vertices points in the fiber and edges $(u,v)$ for all $u-v\in\pm\mathcal{B}$. Given a fixed $A$, a set $\mathcal{B}$ is called a Markov basis of $A$ if this fiber graph is connected for all vectors $b$. That such a set exists and is finite is a consequence of the Hilbert Basis Theorem, invoking the Fundamental Theorem of Markov bases [12].
@@
Any proper subset of a Markov basis is likely to leave at least one fiber disconnected for a given $A$. The easiest subset to compute, which is at least needed to generate the lattice of correct dimension, is the lattice basis of $A$, which is simply the linear-algebra basis of the vector space $\ker_{\mathbb{Z}}(A)$. It is known that lattice bases leave fibers disconnected in general. On the other hand, work as early as [9] proves that lattice bases suffice for sampling a relaxation of the fiber in the case of logistic regression models; this idea was also successfully explored for no-3-factor models for small tables by [39]. However, the negative results in [3] warn against the generalization of this idea: one has to choose the cells allowing to go negative very carefully, otherwise the chain remains disconnected for arbitrary relaxations.
@@ 2 Multilevel Monte Carlo Methods
Multilevel Monte Carlo (MLMC) methods [17] reduce this cost by using a hierarchy of approximations to the quantity of interest. Rather than estimating $\mathbb{E}[g(\mathbf{X})]$ directly at high resolution, MLMC expresses it as a telescoping sum over a sequence of increasingly accurate (and expensive) models:
@@
This cost saving regime has therefore been applied to many other domains, e.g., [24, 37] despite originally developed for estimating expectations in numerical simulation (e.g., for path dependent finance problems and PDEs). MLMC has more recently been extended to the approximation of distribution functions and densities in [18]. These MLMC generalizations rely on smoothing techniques more sophisticated than standard kernel density estimation. Estimating distribution functions or densities from multilevel procedures directly involves discontinuities or singularities (from e.g., subtraction of indicator functions or Dirac deltas respectively), which can lead to high variance between levels and bias. By replacing these with smooth approximations, MLMC maintains variance decay and ensures stable convergence; see Section 3.1 and 3.2 for full details.
@@ 3 A Multilevel Algorithm for Fiber Sampling
Traditionally, the most used method for sampling fibers has been Metropolis-Hastings (MH) algorithm and its variations. With the introduction of Markov basis [12], such class of algorithms seemed like an obvious and cheap way to obtain samples on the high dimensional lattice bounded by a polytope $\mathcal{P}$. Nevertheless, these algorithms operate on a single sampling level, and their goal is not to efficiently explore the fiber but simply to obtain a sample which follows the proposed distribution. The closest attempt at multilevel fiber sampling was proposed as a lemma in [12].
Lemma 3.1 (Lemma 2, [12]).
@@
The sampling procedure described in Lemma 3.1 defines a valid Markov chain Monte Carlo method, albeit one with acceptance probability $1$. By sampling the step size $j$ from a weighted distribution tailored to the target (hypergeometric) distribution, the algorithm avoids explicit accept/reject steps while preserving reversibility and ensuring the correct stationary distribution. Despite these theoretical guarantees, this sampling procedure like other MH-type methods still requires a burn-in period and long chain lengths to ensure reliable convergence. To illustrate this drawback in practice (see Appendix A for details of the fiber used), we compare the smoothed empirical distributions of chi-squared statistics obtained from two samplers: 1) the original sampler described in Lemma 3.1 and 2) a modified version in which the step size $j\in\mathbb{Z}$ is chosen with probability proportional to
@@
Notice that for the original sampler, most of the sample is concentrated in a narrow region. While one might initially interpret this concentration as statistically meaningful, the ‘inverted’ sampler reveals a drastically different picture when the distribution of step sizes is inverted. In the original sampler, smaller step sizes $j$ are favored, which causes the chain to linger near its starting point. Conversely, when larger $j$ values are prioritized by inverting the probabilities, the chain initially explores more widely but eventually gets trapped near the extremal points of the fiber, again failing to explore its full structure. These observations underscore a central limitation: without careful tuning, naive MH samplers may fail to produce representative samples of the fiber.
To address this, we now propose a multilevel sampling algorithm, adapted to estimate densities from samples generated over a discrete space, that avoids unnecessary burn-in while still producing samples informative for tasks such as goodness-of-fit testing. Our approach follows [18] and combines ideas from both Metropolis–Hastings and multilevel Monte Carlo (MLMC) methods. We focus on fibers which, although a relative term, could be called dense and use moves coming from either Markov basis or lattice basis.
@@ 3.1 Our Multilevel Algorithm
To control variance and ensure stable density estimation, we apply a smoothing kernel $g^{k,\delta}$ to the samples. The kernel function $g$ is typically a compactly supported, $r$-smooth function (e.g., a Matérn kernel), and the evaluation is performed at a fixed set of grid points $s_{1},\ldots,s_{k}$. Specifically, for a sample value $t$, we define
@@
[…]
ensuring unbiasedness of the estimator. That is, the marginal distributions of $Y^{(l)}$ and $Y^{(l-1)}$ must be preserved, and their joint distribution must match across adjacent levels. In practice, we approximate this condition by sampling $Z_{i}^{(l)}$ uniformly at random from the set of samples obtained at level $l-1$. While this approach preserves the correct marginals, it does not enforce exact joint dependence across levels. Nevertheless, it yields an unbiased estimator, and more sophisticated coupling strategies could be employed in future work to reduce variance.
@@
A full algorithmic implementation of our multilevel estimator is given in Appendix B.
@@ 3.2 Theoretical Guarantees
The theoretical foundation for employing the multilevel density estimator (2) follows from the analysis of [18]. The goal is to approximate the true density $\rho$ of a real-valued random variable $Y$ at a fixed set of evaluation points $s_{1},\ldots,s_{k}$. A smoothing kernel $g^{k,\delta}$ regularizes each sample, ensuring that the resulting density estimate is sufficiently smooth for analysis. Crucially, the approximation quality can be controlled by tuning parameters such as the kernel bandwidth $\delta$, the number of grid points $k$, and the sample allocation $N_{\ell}$ across levels to balance bias and variance.
Under mild assumptions on the smoothness of $\rho$, the convergence rates of the level-wise approximations, and the regularity of the kernel $g$, the MSE satisfies the following upper bound (see Equation 3.6 of [18]):
[…]
@@
[…]
@@ 4 Measuring Sample Quality
The problem of comparing probability distributions arises across many areas of modern science and engineering. For example, in quasi-Monte Carlo (QMC) methods, a key task is to “closely” approximate a target distribution by a discrete set of sampling nodes [25, 13]. The quality of such approximations is traditionally measured by discrepancy, often with respect to the uniform distribution on the $d$-dimensional unit hypercube; a setting intrinsic to classical QMC methods.
By contrast, in algebraic statistics, where samples are drawn from fibers defined by algebraic constraints, there is no analog of a discrepancy metric to quantify how well a finite sample represents the underlying space. While MCMC practitioners routinely use convergence diagnostics and appeal to concepts like exchangeability [8, 10], these provide only indirect assurance of sample quality. A sample may be exchangeable yet still fail to meaningfully capture the structure of the fiber. Motivated by this gap, we consider metrics for comparing discrete distributions, or for evaluating how thoroughly a single sample explores its fiber.

### v2

@@ 1 Introduction
Henry Wynn’s work in bringing modern algebraic geometry to the field of statistics laid the foundation for the development of algebraic statistics as a subject. The term ‘algebraic statistics’ was coined in his joint work [PistoneWynnEva] (see also the monograph [WynnAlgStatBook2001]) and broadly denotes the application of nonlinear algebraic and nonconvex geometric methods to statistical problems. Although algebraic and geometric ideas had long been present in statistics, the predominant tools before the 2000s were linear algebra and convex geometry. A shift began in the mid-1990s: Pistone and Wynn [PistoneWynn] employed Gröbner bases to study confounding in experimental design, while Diaconis and Sturmfels [DS98] used them for performing exact conditional tests. Wynn and collaborators went further, recasting core concepts in statistics–such as optimal design and system reliability theory [reliability]–within the framework of computational algebra. Today, Wynn’s influence remains central in these areas. A central theme of his work was bridging interdisciplinary language barriers to connect ideas across fields. This paper presents one such effort [^1] In 2016, Henry Wynn visited Chicago twice. During discussions with S. Petrović, F. Hickernell, and their students, he proposed a new approach to sampling lattice points in polytopes, inspired by multilevel methods.: the adaptation of multilevel sampling [Gil15a]–a powerful tool in Monte Carlo methods employed to reduce computational cost when approximating expectations of quantities of interest, most commonly applied to path simulation [Gil08a]–to the setting of sampling discrete structures, a core concern in algebraic statistics.
@@
Sampling algorithms and strategies for exploring the set (1) abound. These include, for example, the hit-and-run sampler [hit-and-run] (see also [DiaconisAnderson]), sequential importance sampling (SIS, see [Liu2001reprint] as a standard reference), and a hybrid approach that combines SIS and Markov bases [KahleYoshidaGarciaPuente]. [Chen_etal-lattice-contingency] reviews and compares a Markov chain implementation and SIS for logistic regression models, showing how SIS outperforms in certain scenarios; while [Dob2012] demonstrates the superiority of Markov bases over SIS in general fibers. Dobra also focused on the problem of actually computing Markov bases in practice, which is intractable for many problems of reasonable size that arise in applications. Indeed, there have been many attempts to work without the entire basis; see [HAT] for an early idea based on remarks made in [DS98]. The fact of the matter is that using a proper subset of a Markov basis necessarily disconnects some fibers, and it is impossible to tell in general which ones. A recent approach to circumvent the problem is a Bayesian framework, called RUMBA [RUMBA], that learns combinations of smaller basis elements. Given recent advances in reinforcement learning, another reasonable way to try to solve the problem is by actor-critic algorithms; this has been proposed in [RLFiberSampler], motivated by [DeLoeraWu-RL]. Today, the fiber sampling problem still faces some open challenges: training agents comes at a cost, tuning sampler parameters takes time, and mixing of Markov chains is difficult to evaluate in a timely manner.
@@ 1.1 Problem and notation
Most of our readers will be familiar with the fact that conditional distributions given sufficient statistics in log-affine models, as defined in [Lauritzen], are supported exactly on lattice points in polytopes. Log-affine models are exponential family models ([Barndorff-Nielsen]) for $k$ discrete random variables $X_{1},\dots,X_{k}$ with finite state spaces, $X_{i}\subset[d_{i}]:=\{1,\dots,d_{i}\}$. An exponential family being log-affine implies the existence of a large, often sparse matrix $A=[a_{1}\dots a_{n}]\subset\mathbb{Z}^{d\times n}$, which computes the model’s sufficient statistics: the family of probability distributions $\mathcal{M}_{A}:=\{p_{\theta}\}_{\theta\in\Theta}$ takes the form:
@@
One canonical application for constructing a sample from (1) is to perform the test of goodness-of-fit for $\mathcal{M}_{A}$ when asymptotic considerations fail–namely, in high-dimensional or large but possibly sparse state spaces. Early references such as [haberman1976generalized, Habe:1981, Haberman77, Haberman88] include discussions inadequacy of standard asymptotic approximations are for estimating the model fit $p$-value based on the approximate distribution of the chi-squared statistic. Thus, in our demonstrations, we focus obtaining the sampling distribution of the chi-square statistic, as it can be used as a goodness-of-fit statistic for testing the fit of the model to the observed data. The distribution on the fiber can be either uniform (for conditional volume tests [DiaconisEfron-volumeTest]) or the conditional distribution given sufficient statistics (uniform or often hypergeometric, depending on the table cell sampling scheme [FienberBookCategorical]). Measuring distance of observed data to the set of joint observations consistent with the log-linear model naturally lead to conditioning on sufficient statistics, removing dependence on model parameters; conditional inference was described by Agresti in 1990 (see [Agresti2002, Agresti]). This was the context considered in [DS98], who introduced Markov bases for solving this broad class of sampling problems. The crucial fact is that a Markov chain is built using moves that have analogues in polynomial algebra and polyhedral geometry. Each instance inherits a finite set of moves necessary to obtain an irreducible chain for any value of the sufficient statistic. Diaconis and Sturmfels named any such finite collection of moves a Markov basis; see [DS98, Theorem 3.1]. Their seminal paper has over 1000 citations; we point the readers to the comprehensive references [AHT2012] and [SethBook] and a recent survey [MB25years].
@@
To fix the notation, we will refer to the set $\mathcal{F}_{A}(b)$} from Equation (1) as the $b$-fiber of $A$. Elements ${\mathcal{B}}\subset\ker_{\mathbb{Z}}(A)$, which by definition preserve the value of the sufficient statistics, are called moves on the fiber. For given moves $\mathcal{B}$, one can construct a fiber graph with vertices points in the fiber and edges $(u,v)$ for all $u-v\in\pm\mathcal{B}$. Given a fixed $A$, a set $\mathcal{B}$ is called a Markov basis of $A$ if this fiber graph is connected for all vectors $b$. That such a set exists and is finite is a consequence of the Hilbert Basis Theorem, invoking the Fundamental Theorem of Markov bases [DS98].
@@
Any proper subset of a Markov basis is likely to leave at least one fiber disconnected for a given $A$. The easiest subset to compute, which is at least needed to generate the lattice of correct dimension, is the lattice basis of $A$, which is simply the linear-algebra basis of the vector space $\ker_{\mathbb{Z}}(A)$. It is known that lattice bases leave fibers disconnected in general. On the other hand, work as early as [Chen_etal-lattice-contingency] proves that lattice bases suffice for sampling a relaxation of the fiber in the case of logistic regression models; this idea was also successfully explored for no-3-factor models for small tables by [Yoshida-Barnhill:Negative-cell]. However, the negative results in [MB25years] warn against the generalization of this idea: one has to choose the cells allowing to go negative very carefully, otherwise the chain remains disconnected for arbitrary relaxations.
@@ 2 Multilevel Monte Carlo Methods
Multilevel Monte Carlo (MLMC) methods [Gil15a] reduce this cost by using a hierarchy of approximations to the quantity of interest. Rather than estimating $\mathbb{E}[g(\mathbf{X})]$ directly at high resolution, MLMC expresses it as a telescoping sum over a sequence of increasingly accurate (and expensive) models:
@@
This cost saving regime has therefore been applied to many other domains, e.g., [HerSch20a, Vro19a] despite originally developed for estimating expectations in numerical simulation (e.g., for path dependent finance problems and PDEs). MLMC has more recently been extended to the approximation of distribution functions and densities in [gil15_multi_density]. These MLMC generalizations rely on smoothing techniques more sophisticated than standard kernel density estimation. Estimating distribution functions or densities from multilevel procedures directly involves discontinuities or singularities (from e.g., subtraction of indicator functions or Dirac deltas respectively), which can lead to high variance between levels and bias. By replacing these with smooth approximations, MLMC maintains variance decay and ensures stable convergence; see Section 3.1 and 3.2 for full details.
@@ 3 A Multilevel Algorithm for Fiber Sampling
Traditionally, the most used method for sampling fibers has been Metropolis-Hastings (MH) algorithm and its variations. With the introduction of Markov basis [DS98], such class of algorithms seemed like an obvious and cheap way to obtain samples on the high dimensional lattice bounded by a polytope $\mathcal{P}$. Nevertheless, these algorithms operate on a single sampling level, and their goal is not to efficiently explore the fiber but simply to obtain a sample which follows the proposed distribution. The closest attempt at multilevel fiber sampling was proposed as a lemma in [DS98].
Lemma 3.1 (Lemma 2.2, [DS98]).
@@
The sampling procedure described in Lemma 3.1 defines a valid Markov chain Monte Carlo method, albeit one with acceptance probability $1$. By sampling the step size $j$ from a weighted distribution tailored to the target (hypergeometric) distribution, the algorithm avoids explicit accept/reject steps while preserving reversibility and ensuring the correct stationary distribution. The authors further illustrate how a different stationary distribution on the fiber can be reached by this Markov chain by changing the probability of choosing the $j$s. Despite these theoretical guarantees, this sampling procedure like other MH-type methods still requires a burn-in period and long chain lengths to ensure reliable convergence.
One may be tempted to use this result as a basis for a multilevel sampler. However, this method is more like a Gibbs sampler (or hit-and-run), rather than a multilevel sampler. In fact, attempting to control the sizes of moves in the Markov chain by changing the multiples $j$ for any particular Markov move using this result looses control of the stationary distribution. To illustrate this issue (see Appendix A for details of the fiber used), we compare the smoothed empirical distributions of chi-squared statistics obtained from two samplers: 1) the original sampler described in Lemma 3.1 and 2) a modified version in which the step size $j\in\mathbb{Z}$ is chosen with probability proportional to
@@
The original sampler’s stationary distribution of the chi-square statistic is induced by the hypergeometric distribution of the fiber elements, as per Lemma 3.1. Notice that smaller step sizes $j$ are favored, which causes the chain to linger near its starting point. The second sampler prioritizes larger step sizes $j$ by inverting the probabilities; however, the stationary distribution is, by design, induced by another distribution on the fiber. These observations underscore a central limitation: without careful tuning of step sizes, naive MH samplers may fail to produce representative samples of the fiber, or may converge to a different stationary distribution.
To address this, we now propose a multilevel sampling algorithm, adapted to estimate densities from samples generated over a discrete space, that avoids unnecessary burn-in while still producing samples informative for tasks such as goodness-of-fit testing. Our approach follows [gil15_multi_density] and combines ideas from both Metropolis–Hastings and multilevel Monte Carlo (MLMC) methods. We focus on fibers which, although a relative term, could be called dense and use moves coming from either Markov basis or lattice basis.
@@ 3.1 Our Multilevel Algorithm
To control variance and ensure stable density estimation, we apply a smoothing kernel $g^{k,\delta}$ to the samples. The kernel function $g$ is typically a compactly supported, $r$-smooth function and the evaluation is performed at a fixed set of grid points $s_{1},\ldots,s_{k}$. Specifically, for a sample value $t$, we define
@@
$(Y_{i}^{(l)},Z_{i}^{(l)})\overset{d}{=}(Y^{(l)},Y^{(l-1)}),$
ensuring unbiasedness of the estimator. That is, the marginal distributions of $Y^{(l)}$ and $Y^{(l-1)}$ must be preserved, and their joint distribution must match across adjacent levels. In practice, we approximate this condition by sampling $Z_{i}^{(l)}$ uniformly at random from the set of samples obtained at level l1l-1. While this approach preserves the correct marginals, it does not enforce exact joint dependence across levels. Nevertheless, it yields an unbiased estimator, and more sophisticated coupling strategies could be employed in future work to reduce variance.
@@
A full algorithmic implementation of our multilevel estimator is given in Appendix B. In our implementation, we use the conditional volume test and generate uniform samples from the fiber by accepting all feasible Markov moves. Alternative target distributions on the fiber can be obtained by redefining the acceptance probability of the Markov chain when generating the $N_{l}$ samples in Algorithm 1.
@@ 3.2 Theoretical Guarantees
The theoretical foundation for employing the multilevel density estimator (2) follows from the analysis of [gil15_multi_density]. The goal is to approximate the true density $\rho$ of a real-valued random variable $Y$ at a fixed set of evaluation points $s_{1},\ldots,s_{k}$. A smoothing kernel $g^{k,\delta}$ regularizes each sample, ensuring that the resulting density estimate is sufficiently smooth for analysis. Crucially, the approximation quality can be controlled by tuning parameters such as the kernel bandwidth $\delta$, the number of grid points $k$, and the sample allocation $N_{\ell}$ across levels to balance bias and variance.
Under mild assumptions on the smoothness of $\rho$, the convergence rates of the level-wise approximations, and the regularity of the kernel $g$, the MSE satisfies the following upper bound (see Equation 3.6 of [gil15_multi_density]):
$\text{MSE}\lesssim\$$k^{-2r}+\delta^{2r}$
@@
$+\frac{\log k}{\delta^{2}}\cdot\left(\frac{1}{N_{l_{L}}}+\sum_{\ell=l_{1}}^{l_{L}-1}\frac{\min\left(\delta^{-\beta_{4}}M^{-\beta_{5}\ell},\ 1\right)}{N_{\ell}}\right)$
@@ 4 Measuring Sample Quality
The problem of comparing probability distributions arises across many areas of modern science and engineering. For example, in quasi-Monte Carlo (QMC) methods, a key task is to “closely” approximate a target distribution by a discrete set of sampling nodes [HICKKIRKSOR2025, DicEtal14a]. The quality of such approximations is traditionally measured by discrepancy, often with respect to the uniform distribution on the $d$-dimensional unit hypercube; a setting intrinsic to classical QMC methods.
By contrast, in algebraic statistics, where samples are drawn from fibers defined by algebraic constraints, there is no analog of a discrepancy metric to quantify how well a finite sample represents the underlying space. While MCMC practitioners routinely use convergence diagnostics and appeal to concepts like exchangeability [BesagClifford89, Diaconis2022exchangabilityTables], these provide only indirect assurance of sample quality. A sample may be exchangeable yet still fail to meaningfully capture the structure of the fiber. Motivated by this gap, we consider metrics for comparing discrete distributions, or for evaluating how thoroughly a single sample explores its fiber.